OpenStudy (anonymous):

(will award medal!!) check my work??? To travel 60 miles, it takes sue, riding a Harley, 2 hours less time than it takes Matt to travel 50 miles riding a bicycle. Sue travels 10 miles per hour faster than Matt. Find the times and rates of Matt and sue.

OpenStudy (anonymous):

ahh someone please help? I'm new to this site and this question is hard for me :/

OpenStudy (anonymous):

I already have SR= 60/ST ST= MT-2 MR=50/MT SR=MR+10

OpenStudy (anonymous):

just want to know if I'm right!?

OpenStudy (dumbcow):

yes the equations are correct

OpenStudy (dumbcow):

use substitution to get equation with 1 variable....then work backwards to get other variables

OpenStudy (anonymous):

wait so that's not the final answer?

OpenStudy (dumbcow):

--> 60/ST = MR + 10 --> 60/(MT -2) = MR +10 --> 60/(50/MR -2) = MR +10

OpenStudy (anonymous):

whats that?

OpenStudy (dumbcow):

its using substitution to solve for MR

OpenStudy (dumbcow):

you have 4 variables and 4 equations you have to solve the system

OpenStudy (anonymous):

I'm confused... I didn't know I needed to solve for MR? i thought i just had to find the rates and times of "Matt and Sue" .. the questions doesnt ask me to solve?

OpenStudy (dumbcow):

its the same thing :) you need numbers for how fast each are going and how long it takes them

OpenStudy (anonymous):

oh okay... so if I'm using the subsitution method which one do I solve first?

OpenStudy (dumbcow):

doesn't matter....just get an equation with only 1 variable

OpenStudy (anonymous):

like taking SR=MR+10 and making it 10(SR)=MR ??

OpenStudy (anonymous):

idk if thats not right im really bad at this... whoever can just post all the work to solve this question and then explain it to help me understand it from there? I'm usually better at understanding if I can see the question and the work at the same time..

OpenStudy (dumbcow):

ok im not sure what you just did there ^^ have you solved a system of 2 equations before?

OpenStudy (anonymous):

yes but a really long time ago..

OpenStudy (anonymous):

im not really sure where to even start with this one

OpenStudy (dumbcow):

SR = MR + 10 --> 60/ST = MR + 10 (SR = 60/ST) --> 60/(MT -2) = MR +10 (ST = MT-2) --> 60/(50/MR -2) = MR +10 (MT = 50/MR) does this make sense?

OpenStudy (dumbcow):

from here you can simplify and solve for MR \[\frac{60}{\frac{50-2MR}{MR}} = MR +10\] \[60MR = (MR +10)(50 -2MR)\] \[2MR^2 + 30MR - 500 = 0\] Factor \[2(MR +25)(MR -10) = 0\] MR must be positive \[MR = 10\]

OpenStudy (whpalmer4):

Ssue = 60/Tsue Smatt = 50/Tmatt Ssue = Smatt + 10 Tsue = Tmatt - 2 Substituting in some of the relationships to get everything in terms of Tmatt: Smatt+10 = 60/(Tmatt - 2) 10 + 50/Tmatt = 60/(Tmatt-2) (10 Tmatt + 50)/Tmatt = 60/(Tmatt-2) (Tmatt-2)(10 Tmatt + 50) = 60 Tmatt 10 Tmatt^2 + 50 Tmatt - 20 Tmatt - 100 = 60 Tmatt 10 Tmatt^2 - 30 Tmatt - 100 = 0 Divide by 10 to make numbers smaller Tmatt^2 - 3 Tmatt - 10 = 0 Solve by factoring or quadratic Tmatt = -2, Tmatt = 5 Tmatt = 5 is the only solution that makes sense in the context of this problem. From there you can find the others.